what is another equation for y=5x-3
Answer by KMST(5328)

There are infinity of expressions representing the same line.
y=5x-3 could be said to be the equation of the line in the slope-intercept form, that shows y as a function of x, with the slope of the line (3) multiplying the x and the y for the y-intercept point as the independent term (-3).
You could express the same line in the standard form
5x-y-3=0 or 5x-y=3.
You could also multiply both sides of any equation by the same non-zero number to get an equivalent expression. In particular
x/(3/5) + y/(-3) = 1
shows you the x-intercept and y-intercept dividing x and y respectively.